# A Practical Validation of RIS Detection and Identification

### v1

@@ IV-A RIS 1 Reachable Scenario
In this scenario, all reflecting elements are assigned to RIS 1. Accordingly, the states of elements are adjusted according to the ARP of RIS 1. As seen in Fig. 4, when only RIS 1 is reachable, $\overline{D^{(2)}_{\text{max}}}$ has low value and does not increase with the increase in Tx gain. On the other hand, $\overline{D^{(1)}_{\text{max}}}$ is consistently higher than $\overline{D^{(2)}_{\text{max}}}$, and increases with the Tx gain. For example, increasing the Tx gain by $3$ dB approximately doubles $\overline{D^{(1)}_{\text{max}}}$. Additionally, in Fig. 5, the effect of the ARP length $M$ on $\overline{D^{(1)}_{\text{max}}}$ is investigated. It can be seen that $\overline{D^{(1)}_{\text{max}}}$ increases with the increase of $M$ where doubling $M$ increases the $\overline{D^{(1)}_{\text{max}}}$ by a factor of two, due to the increase in the correlation amplitude.
In Fig. 6, the effect of the Tx gain on $P_{miss}$ of RIS 1 and $P_{F}$ of RIS 2 is examined, with the x-axis of the figure representing the normalized detection threshold $\widetilde{thr}$. Here, $\widetilde{thr}=thr/P_{noise}$, where $P_{noise}$ is the noise power measured at the Rx SDR. From Fig. 6, it is observed that while increasing the Tx gain decreases $P_{miss}$ of RIS 1, it does not affect $P_{F}$ of RIS 2 considerably. Additionally, it is clearly seen that increasing $\widetilde{thr}$ decreases $P_{F}$ of RIS 2 while increases $P_{miss}$ of RIS 1 as expected [9].
@@ IV-B Both RISs Reachable Scenario
Figure 7: The effect of the Tx gain on $\overline{D^{(l)}_{\text{max}}}$ for RIS 1 and RIS 2 in ”Both RISs reachable” scenario.
In this scenario, while approximately half of reflecting elements are assigned to RIS 1, the other half are assigned to RIS 2 as seen in Fig. 3. Accordingly, the states of elements are adjusted based on their assignments: those assigned to RIS 1 are configured according to the ARP of RIS 1, while those assigned to RIS 2 are configured according to the ARP of RIS 2. Fig. 7 shows the effect of the Tx gain on $\overline{D^{(l)}_{\text{max}}}$ for RIS 1 and RIS 2 when both RISs are reachable. The experiment is performed for two different positions of the UE. It is observed that while $\overline{D^{(1)}_{\text{max}}}$ in position 1 compared to $\overline{D^{(2)}_{\text{max}}}$, the opposite is true in position 2. This behaviour is resulted from that the transmitted signal can not be focused on both RISs fairly due to the miss alignment between the different components of the experiment. Additionally, it is seen that increasing the Tx gain increases $\overline{D^{(l)}_{\text{max}}}$ for both RISs where increasing the Tx gain $3$ dB leads to and approximately two-fold increase in $\overline{D^{(l)}_{\text{max}}}$ values. In Fig. 8, we show the impact of $M$ on $\overline{D^{(l)}_{\text{max}}}$ for RIS 1 and RIS 2 when both RISs are reachable. It is seen that increasing $M$ increases $\overline{D^{(l)}_{\text{max}}}$ for both RISs, where doubling $M$ doubles $\overline{D^{(l)}_{\text{max}}}$ for both RISs.
In Fig. 9, the effect of $\widetilde{thr}$ on $P_{miss}$ of RIS 1 and RIS 2 is investigated. The results show that increasing $\widetilde{thr}$ leads to a higher $P_{miss}$ for both RISs. Specifically, when the Tx gain is $-50$ dB, increasing $\widetilde{thr}$ from $0.6$ to $0.8$ increases $P_{miss}$ from $0.23$ to $0.31$ for RIS 1 and from $0.09$ to $0.22$ for RIS 2. Moreover, increasing the Tx gain decreases the $P_{miss}$ for both RISs. This behavior is evident from Fig. 9 for RIS 2. For example, when $\widetilde{thr}$ is set at $0.8$, increasing the Tx gain by $10$ dB reduces the $P_{miss}$ for RIS 1 from $0.22$ to $0.013$.
@@ IV-C No RIS Reachable Scenario
In the scenario, RIS is removed from the environment, and identification process is performed using Algorithm 1. Then, the effect of $\widetilde{thr}$ on $P_{F}$ of RIS 1 and RIS 2 is investigated. As seen in Fig. 10, increasing $\widetilde{thr}$ decreased $P_{F}$ of RIS 1 and RIS 2. For example, when the Tx gain is $-50$ dB, increasing $\widetilde{thr}$ from $1.4$ to $1.8$, decreased $P_{F}$ of RIS 1 from $0.67$ to $0.23$ and decreased $P_{F}$ of RIS 2 from $0.27$ to $0.1$.

### v2

@@ IV-A RIS 1 Reachable Scenario
In this scenario, all reflecting elements are assigned to RIS 1. Accordingly, the states of elements are adjusted according to the ARP of RIS 1. As seen in Fig. 3, when only RIS 1 is reachable, $\overline{D^{(2)}_{\text{max}}}$ has low value and does not increase with the increase in Tx gain or ARP length $M$. On the other hand, $\overline{D^{(1)}_{\text{max}}}$ is consistently higher than $\overline{D^{(2)}_{\text{max}}}$, and increases with both the Tx gain and $M$. For example, increasing the Tx gain by $3$ dB approximately doubles $\overline{D^{(1)}_{\text{max}}}$ and, doubling $M$ increases the $\overline{D^{(1)}_{\text{max}}}$ by a factor of two, due to the increase in the correlation amplitude.
In Fig. 4, the effect of the Tx gain on $P_{miss}$ of RIS 1 and $P_{F}$ of RIS 2 is examined, with the x-axis of the figure representing the normalized detection threshold $\widetilde{thr}$. Here, $\widetilde{thr}=thr/P_{noise}$, where $P_{noise}$ is the noise power measured at the Rx SDR. From Fig. 4, it is observed that while increasing the Tx gain decreases $P_{miss}$ of RIS 1, it does not affect $P_{F}$ of RIS 2 considerably. Additionally, it is clearly seen that increasing $\widetilde{thr}$ decreases $P_{F}$ of RIS 2 while increases $P_{miss}$ of RIS 1 as expected [RIS_ID_Theoretical].
@@ IV-B Both RISs Reachable Scenario
In this scenario, while approximately half of reflecting elements are assigned to RIS 1, the other half are assigned to RIS 2. Accordingly, the states of elements are adjusted based on their assignments: those assigned to RIS 1 are configured according to the ARP of RIS 1, while those assigned to RIS 2 are configured according to the ARP of RIS 2. In Fig. 5, we show the effect of $M$ and Tx gain on $\overline{D^{(l)}_{\text{max}}}$ for RIS 1 and RIS 2 when both RISs are reachable. It is observed that $\overline{D^{(1)}_{\text{max}}}$ is higher than $\overline{D^{(2)}_{\text{max}}}$. This behavior is resulted from that the transmitted signal can not be focused on both RISs fairly due to the miss alignment between the different components of the experiment. Additionally, it is seen that increasing the Tx gain increases $\overline{D^{(l)}_{\text{max}}}$ for both RISs where increasing the Tx gain $3$ dB leads to approximately two-fold increase in $\overline{D^{(l)}_{\text{max}}}$ values. Similarly, increasing $M$ also increases $\overline{D^{(l)}_{\text{max}}}$ for both RISs, where doubling $M$ doubles $\overline{D^{(l)}_{\text{max}}}$.
In Fig. 6, the effect of $\widetilde{thr}$ on $P_{miss}$ of RIS 1 and RIS 2 is investigated. The results show that increasing $\widetilde{thr}$ leads to a higher $P_{miss}$ for both RISs. Specifically, when the Tx gain is 50-50 dB, increasing $\widetilde{thr}$ from $0.4$ to $1.2$ increases $P_{miss}$ from $0.12$ to $0.41$ for RIS 1 and from $0.03$ to $0.53$ for RIS 2. Moreover, increasing the Tx gain decreases the $P_{miss}$ for both RISs. This behavior is evident from Fig. 6 for RIS 2. For example, when $\widetilde{thr}$ is set at $1$, increasing the Tx gain by $10$ dB reduces the $P_{miss}$ for RIS 2 from $0.35$ to $0.03$.
@@ IV-C No RIS Reachable Scenario
In this scenario, the RIS is removed from the environment, and the identification process is performed using Algorithm 1. Then, the effect of $\widetilde{thr}$ on $P_{F}$ of RIS 1 and RIS 2 is investigated. As seen in Fig. 7, increasing $\widetilde{thr}$ decreased $P_{F}$ of RIS 1 and RIS 2. For example, when the Tx gain is 50-50 dB, increasing $\widetilde{thr}$ from $1$ to $1.6$, decreased $P_{F}$ of RIS 1 from $0.25$ to $0.06$ and decreased $P_{F}$ of RIS 2 from $0.63$ to $0.20$.
Finally, we note that although the effect of the overall Tx–RIS–Rx link distance on the performance metrics was not explicitly investigated in any scenarios of the experiment, it can be inferred from the results in which Tx gain was varied. Increasing the Tx gain enhances the signal-to-noise ratio (SNR) at the receiver and thereby increases $\overline{D^{(l)}_{\text{max}}}$ for reachable RISs, as shown in Figs. 3 and 5. This increase makes it more likely for $\overline{D^{(l)}_{\text{max}}}$ to exceed the detection threshold for reachable RISs. As a result, a higher Tx gain leads to a lower $P_{miss}$, while the $P_{F}$ remains mostly unchanged, as observed in Figs. 4, 6, and 7. On the other hand, increasing the Tx–RIS–Rx distance would introduce higher path loss, which in turn reduces the SNR at the receiver and consequently decreases $\overline{D^{(l)}_{\text{max}}}$ for reachable RISs. Therefore, increasing the Tx–RIS–Rx distance is analogous to decreasing the Tx gain and would be expected to increase $P_{miss}$ of reachable RISs while not changing $P_{F}$ of unreachable RISs considerably.